OpenStudy (anonymous):

how do is solve this trig limit [1-tanx]/sinx-cosx

OpenStudy (jamesj):

as x -> what?

OpenStudy (anonymous):

as x aproaches pi/4

OpenStudy (jamesj):

Note that 1 - tan x = 1/cos x . ( cos x - sin x) Hence the ratio (1 - tan x)/(sin x - cos x) = -1/cos x Now you can take the limit.

OpenStudy (anonymous):

ooh thanks